# Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach

### v3

@@ Abstract
where the total transition rate is $k_{\text{tot}}=r+\gamma_{0}+\gamma_{1}$. This choice fully eliminates the fastest decaying relaxation mode. In addition, for a sufficiently large reset rate $r$, the slowest mode’s amplitude becomes exponentially small after the reset phase. Alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential. Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ serves as a useful $\rho_{\delta}$ regardless of the rate $r$. Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state decelerates relaxation, necessitating the choice $\rho_{\delta}=|1\rangle\langle 1|$ for acceleration. This example demonstrates that, by appropriately choosing $\rho_{\delta}$, one can switch between reset-induced acceleration and deceleration of relaxation.
@@
We next present numerical results for illustration. We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{\rm ss}$ using the standard trace distance $D[\rho(t)|\rho_{\rm ss}]:=\text{Tr}|\rho(t)-\rho_{\rm ss}|/2$ where $|A|:=\sqrt{A^{\dagger}A}$. Other measures, such as the $L_{\infty}$ norm $D_{\infty}[\rho(t)|\rho_{\rm ss}]:=\max_{i}|\lambda_{i}|$ [$\lambda_{i}$ is the $i$-th eigenvalue of $\rho(t)-\rho_{\rm ss}$], yield qualitatively similar outcomes. We first plot $D[\rho(t)|\rho_{\rm eq}]$ over $t$ for various rate $r$ in Fig. 1 (a). The parameters for $r>0$ cases are: initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$; The $r=0$ scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$. The environment is at lower temperature $\beta_{\text{env}}=4.0$. Although the $r>0$ cases are initially farther from equilibrium compared to the $r=0$ case, the reset protocol substantially accelerates relaxation so that the former cases reach equilibrium even faster. We next plot $t_{s}$ as a function of $r$ in Fig. 1 (b) and (d), which show that $t_{s}$ decreases monotonically as $r$ increases. Fig. 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the $L_{\infty}$ norm to reduce fluctuations in the curves.
@@
Figure 1: Relaxation dynamics of the two-state system. (a) and (c): Distances between $\rho(t)$ and $\rho_{\text{eq}}$ as the function of time $t$ with different resetting rate $r$. (b) and (d): The $t_{s}$ as a function of the resetting rate $r$. The parameters are chosen as $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$ with reset protocol and $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$ without reset. For (a) and (b) $\Omega=0$, for (c) and (d) $\Omega=2.0$.
@@
and jump operators $J_{i,\downarrow}=\sqrt{\gamma}\,\sigma_{i}^{-},\quad J_{i,\uparrow}=\sqrt{\gamma e^{-\beta}}\;\sigma_{i}^{+}$. Here, $\sigma^{x,z}_{i}$ are Pauli operators, $\sigma_{i}^{-}=|0\rangle\langle 1|$ and $\sigma_{i}^{+}=|1\rangle\langle 0|$. We fix $J=1$, $g=1.2$, $\gamma=0.5$, $\beta=1/k_{B}T=1$ throughout. We choose $\rho_{\delta}=\mathbb{I}/d$ (with $d=2^{N}$). The initial state $\rho_{0}$ is chosen by normalizing $\rho_{\rm ss}+\alpha V_{2}$, with $\rho_{\rm ss}$ being the steady-state and $V_{2}$ being the second normalized eigenvector of the Lindbladian. A larger $\alpha$ implies that $\rho_{0}$ is farther from $\rho_{\rm ss}$. Here, $\rho_{\rm ss}$ is nontrivial, with nonzero entanglement. Parameters: $t_{s}=0.5\,\tau_{2}$ and $t_{s}=0.08\,\tau_{2}$ with $\tau_{2}=1/|\operatorname{Re}\lambda_{2}|$, the relaxation timescale, total duration $T=6\,\tau_{2}$, $\alpha=0.55/0.05$. Notably, $t_{s}$ is chosen arbitrarily, rather than being determined a prior. The accelerated relaxation processes under different reset rate are shown in Fig. 2. We also examine cases with different parameters in [59], where acceleration is consistently observed. In Fig. 2 (b) and (d), genuine Mpemba effects are observed, where the system is first driven away from the steady state by the reset protocol compared to the no-reset case, and then reaches stationarity faster. The existence of such an effect broadens the range of usable reset states.
Some remarks are in order regarding experimental implications. In general, the optimal duration $t_{s}$ of the reset phase cannot be determined a priori from theoretical considerations, as the relevant dynamical details are typically unknown in complex quantum systems. The TFIM example shows how this challenge can be overcome: fine-tuning of $t_{s}$ is not required—any finite value shorter than the native relaxation time leads to substantial speed-up, and $t_{s}$ can be further adjusted a posteriori based on observed dynamics. This robustness, without requiring prior knowledge of the system dynamics and initial states, distinguishes our protocol from previous approaches. Additionally, the reset channel in the example corresponds to a depolarizing channel, which is conceptually straightforward to implement, e.g., by applying an isotropic white-noise field or coupling an infinite-temperature reservoir.
Figure 2: Acceleration of the relaxation of TFIMs. Reset rate $r$ are chosen as $0,1.0,5.0,10.0,20.0$. (a) $N=5$ ($d=32$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$. (b) $N=5$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$. (c) $N=6$ ($d=64$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$. (d) $N=6$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$. The vertical gray lines indicate $t_{s}$ when the reset channel is removed. Inset: Schematic representation of the dissipative TFIM, with each spin coupled to a thermal bath.
@@
This work is partly supported by the Innovation Program for Quantum Science and Technology (2021ZD0303306), MOST(2022YFA1303100).
@@
$\mathcal{D}_{r}(\rho)=\sum_{i=1}^{d}\!\bigl{(}J_{i}^{r}\rho\,J_{i}^{r\dagger}-\tfrac{1}{2}\{J_{i}^{r\dagger}J_{i}^{r},\rho\}\bigr{)}.$ (19)
@@
Realizing the reset protocol with arbitrary state preparation processes.—We now show that any state preparation protocol could approximately realize the reset channel in our scheme via a Trotterization-based construction. Notably, Trotterization has already been implemented experimentally in superconducting quantum circuits [79], IBM Quantum’s hardware [74] and even quantum many-body platform [77] to generate dissipative dynamics, such as the dissipative TFIM. Moreover, tensor network methods offer promising avenues for extending Trotterization to more complex open quantum systems [72]. We first consider the case where the original dynamics is absent [$\mathcal{L}(\rho)=0$] to gain insight. In this case, the generator corresponding to the reset channel simply reads
@@
That is, a reset channel with rate $r=-\ln p/t_{s}$ is effectively equivalent to a state preparation protocol for $\rho_{\delta}$, with success probability $1-p$. $p$ is a classical probability generated beforehand. One simply flips a biased coin (or, equivalently, performs any local measurement that yields the desired classical randomness): with probability $1-p$ the system is driven into $\rho_{\delta}$ by any suitable operations and with the probability $p$ it is left untouched. Since there are no restrictions on the preparation method, once $\rho_{\delta}$ is locally preparable (e.g. a separable state), this operation can be implemented purely with local operations, even for many-body systems. Specifically, when $\rho_{\delta}=\mathbb{I}/d$, Eq. (25) is a depolarizing channel, which can be directly realized by the 2-design [80] without flipping a coin.
@@
to realize the protocol. With this formula, the reset protocol during $[0,t_{s}]$ can be approximated as
@@
Thus, the reset protocol may be achieved experimentally by performing the aforementioned state preparation operations stroboscopically: at each discrete time point $t_{i}=i\delta t,\ (i=1,...,n)$ within $[0,t_{s}]$, one applies the state preparation mapping from $\rho(t_{i})$ to $\rho_{\delta}$ with a given probability $p_{s}$. There are no constraints on the preparation time $\tau_{\rm prep}$ for $\rho_{\delta}$, but on average it adds a cost of $np_{s}\tau_{\rm prep}$ to the tailored relaxation timescale. Therefore, it is desirable to choose a $\rho_{\delta}$ that can be prepared efficiently, e.g. the maximally mixed state in our TFIM example. This state is convenient to prepare: coupling to an infinite-temperature bath can achieve it in $\mathcal{O}(\ln N)$ time [82] with minimal resources in a $N$-body system—typically negligible compared to the system’s intrinsic relaxation time. Local control methods could in principle realize it in constant time, though with resources scaling with $N$.
@@
$||e^{(\mathcal{L}+\mathcal{R})t_{s}}-(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}||\leq\frac{t_{s}^{2}}{2n}||[\mathcal{L,R}]||,$ (29)
where |||||$\cdot$||| | ⋅ | | denotes any norm that is contractive under Lindbladian evolution, such as the trace norm or the diamond norm. This bound holds for arbitrary Lindbladian $\mathcal{L}$ and $\mathcal{R}$. In our specific case,
$[\mathcal{L,R}]\rho:=[\mathcal{LR}-\mathcal{RL}]\rho=r[\mathcal{L}(\rho_{\delta})-\rho_{\delta}].$ (30)
Hence, a large $n$ is not required when $t_{s}\ll 1$, $rt_{s}\ll 1$ or $\rho_{\delta}$ is close to $\rho_{\rm ss}$. A small $t_{s}$ with moderate $r$ offers a practical regime that may yield substantial acceleration and simultaneously reduce the required number of Trotter steps.
@@
which could be interpreted as a partial and local reset channel (only the $z$-axis is affected, and only local one-body jump operators are involoved). As shown in Fig. 3, the relaxation is accelerated significantly.
Figure 3: Dephasing noise induce acceleration of relaxation in a 5-site TFIM. Parameters: (a) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.2\tau_{2}$ (b) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.8\tau_{2}$.
A theoretical argument supporting the finding is provided in [59].

### v4

@@ Abstract
Since any $t_{s}$ can ensure acceleration, we choose $t_{s}$ such that a faster decaying mode is eliminated. Specifically, solving $c_{4}^{r}(t_{s})=0$ gives such a $t_{s}$, whose expression is provided in [59]. For a sufficiently large $r$, the slowest mode’s amplitude becomes exponentially small after the reset phase. Alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential. Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ serves as a useful $\rho_{\delta}$ regardless of the rate $r$. Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state decelerates relaxation, necessitating the choice $\rho_{\delta}=|1\rangle\langle 1|$ for acceleration. This example demonstrates that, by appropriately choosing $\rho_{\delta}$, one can switch between reset-induced acceleration and deceleration of relaxation.
@@
We next present numerical results for illustration. We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{\rm ss}$ using the standard trace distance $D[\rho(t)|\rho_{\rm ss}]:=\text{Tr}|\rho(t)-\rho_{\rm ss}|/2$ where $|A|:=\sqrt{A^{\dagger}A}$. Other measures, such as the $L_{\infty}$ norm $D_{\infty}[\rho(t)|\rho_{\rm ss}]:=\max_{i}|\lambda_{i}|$ $\lambda_{i}$ is the $i$-th eigenvalue of $\rho(t)-\rho_{\rm ss}$], yield qualitatively similar outcomes. We first plot $D[\rho(t)|\rho_{\rm eq}]$ over $t$ for various rates $r$ in Fig. 1 (a). The parameters for $r>0$ cases are chosen so that they are initially farther from equilibrium compared to the $r=0$ case (see the caption of Fig. 1). The reset protocol substantially accelerates relaxation so that the former cases reach stationarity faster even if they are farther from stationarity initially. We next plot $t_{s}$ as a function of $r$ in Fig. 1 (b) and (d), which show that $t_{s}$ decreases monotonically as $r$ increases. Fig. 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the $L_{\infty}$ norm to reduce fluctuations in the curves. Notably, the steady state is not equilibrium when $\Omega>0$.
@@
Figure 1: Relaxation dynamics of the two-state system. (a) and (c): Distances between $\rho(t)$ and $\rho_{\text{eq}}$ as a function of time $t$ with different resetting rates $r$. (b) and (d): The $t_{s}$ as a function of the resetting rate $r$. The parameters are chosen as $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$ with reset protocol and $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$ without reset. The environment is at lower temperature $\beta_{\text{env}}=4.0$. $\gamma_{1}=1.0$. For (a) and (b) $\Omega=0$, for (c) and (d) $\Omega=2.0$.
@@
and jump operators $J_{i,\downarrow}=\sqrt{\gamma}\,\sigma_{i}^{-},\quad J_{i,\uparrow}=\sqrt{\gamma e^{-\beta}}\;\sigma_{i}^{+}$. Here, $\sigma^{x,z}_{i}$ are Pauli operators, $\sigma_{i}^{-}=|0\rangle\langle 1|$ and $\sigma_{i}^{+}=|1\rangle\langle 0|$. We fix $J=1$, $g=1.2$, $\gamma=0.5$, $\beta=1/k_{B}T=1$ throughout. We choose $\rho_{\delta}=\mathbb{I}/d$ (with $d=2^{N}$). The initial state $\rho_{0}$ is chosen by normalizing $\rho_{\rm ss}+\alpha V_{2}$, with $\rho_{\rm ss}$ being the steady-state and $V_{2}$ being the second normalized eigenvector of the Lindbladian. A larger $\alpha$ implies that $\rho_{0}$ is farther from $\rho_{\rm ss}$. Here, $\rho_{\rm ss}$ is nontrivial, with nonzero entanglement. Moreover, as shown in [59], for a large fraction of randomly chosen pure initial states our protocol with $\rho_{\delta}=\mathbb{I}/d$ still yields substantial acceleration, further demonstrating its robustness. Parameters: $t_{s}=0.5\,\tau_{2}$ and $t_{s}=0.08\,\tau_{2}$ with $\tau_{2}=1/|\operatorname{Re}\lambda_{2}|$, the relaxation timescale, total duration $T=6\,\tau_{2}$, $\alpha=0.55/0.05$. Notably, $t_{s}$ is chosen arbitrarily, rather than being determined a priori. The accelerated relaxation processes under different reset rate are shown in Fig. 2. We also examine cases with different parameters in [59], where acceleration is consistently observed. In Fig. 2 (b) and (d), genuine Mpemba effects are observed, where the system is first driven away from the steady state by the reset protocol compared to the no-reset case, and then reaches stationarity faster. The existence of such an effect broadens the range of usable reset states.
We further verify in [59] that the same protocol significantly accelerates relaxation in another many-body setting, the Dicke model, again using $\rho_{\delta}=\mathbb{I}/d$.
Some remarks are in order regarding experimental implications. In practice, the precise optimal duration $t_{s}$ is generally unknown in advance, but as emphasized before, fine-tuning is unnecessary. The TFIM example directly supports this conclusion: any choice of $t_{s}$ shorter than the intrinsic relaxation time already leads to a clear acceleration. Additionally, the reset channel in the example corresponds to a depolarizing channel, which is conceptually straightforward to implement, e.g., by applying an isotropic white-noise field or coupling an infinite-temperature reservoir.
Figure 2: Acceleration of the relaxation of TFIMs. Reset rates $r$ are chosen as $0,1.0,5.0,10.0,20.0$. (a) $N=5$ ($d=32$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$. (b) $N=5$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$. (c) $N=6$ ($d=64$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$. (d) $N=6$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$. The vertical gray lines mark $t_{s}$ at which the reset channel is removed. Inset: Schematic representation of the dissipative TFIM, with each spin coupled to a thermal bath.
@@
To connect our theoretical protocol more directly with experiments, we provide in [59] proposals that relate $r$ to experimentally tunable parameters, offering realistic paths towards implementation.
@@
We are grateful to anonymous referees for their valuable comments, which greatly enhance the quality of this work. This work is partly supported by the Innovation Program for Quantum Science and Technology (2021ZD0303306), MOST(2022YFA1303100).
@@
$\mathcal{D}_{r}(\rho)=\sum_{i,\alpha}\!\bigl(J_{i,\alpha}^{r}\rho\,J_{i,\alpha}^{r\dagger}-\tfrac{1}{2}\{J_{i,\alpha}^{r\dagger}J_{i,\alpha}^{r},\rho\}\bigr).$ (17)
@@
Realizing the reset protocol with arbitrary state preparation processes.—We now show that any state preparation protocol could approximately realize the reset channel in our scheme via a Trotterization-based construction. Notably, Trotterization has already been implemented experimentally in superconducting quantum circuits [79], IBM Quantum’s hardware [74] and even quantum many-body platform [77] to generate dissipative dynamics, such as the dissipative TFIM. Moreover, tensor network methods offer promising avenues for extending Trotterization to more complex open quantum systems [72]. We first consider the case where the original dynamics is absent $\mathcal{L}(\rho)=0$] to gain insight. In this case, the generator corresponding to the reset channel simply reads
@@
That is, a reset channel with rate $r=-\ln p/t_{s}$ is effectively equivalent to a state preparation protocol for $\rho_{\delta}$, with success probability 1p1-p. $p$ is a classical probability generated beforehand. One simply flips a biased coin (or, equivalently, performs any local measurement that yields the desired classical randomness): with probability 1p1-p the system is driven into $\rho_{\delta}$ by any suitable operations and with the probability $p$ it is left untouched. Since there are no restrictions on the preparation method, once $\rho_{\delta}$ is locally preparable (e.g. a separable state), this operation can be implemented purely with local operations, even for many-body systems. Specifically, when $\rho_{\delta}=\mathbb{I}/d$, Eq. (23) is a depolarizing channel, which can be directly realized by a unitary 2-design [59, 80] without flipping a coin.
@@
to realize the protocol. With this formula, the reset protocol during $0,t_{s}]$ can be approximated as
@@
Thus, the reset protocol may be achieved experimentally by performing the aforementioned state preparation operations stroboscopically: at each discrete time point $t_{i}=i\delta t,\ (i=1,...,n)$ within $0,t_{s}]$, one applies the state preparation mapping from $\rho(t_{i})$ to $\rho_{\delta}$ with a given probability $p_{s}$. There are no constraints on the preparation time $\tau_{\rm prep}$ for $\rho_{\delta}$, but on average it adds a cost of $np_{s}\tau_{\rm prep}$ to the tailored relaxation timescale. Therefore, it is desirable to choose a $\rho_{\delta}$ that can be prepared efficiently, e.g. the maximally mixed state in our TFIM example. This state is convenient to prepare: coupling to an infinite-temperature bath can achieve it in $\mathcal{O}(\ln N)$ time [82] with minimal resources in a $N$-body system—typically negligible compared to the system’s intrinsic relaxation time. Local control methods could in principle realize it in constant time, though with resources scaling with $N$.
@@
$e^{(\mathcal{L}+\mathcal{R})t_{s}}-(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}||\leq\frac{t_{s}^{2}}{2n}$||[$\mathcal{L,R}$]$|$, (27)
where $||\cdot||$ denotes any norm that is contractive under Lindbladian evolution, such as the trace norm or the diamond norm. This bound holds for arbitrary Lindbladian $\mathcal{L}$ and $\mathcal{R}$. In our specific case,
$\mathcal{L,R}]\rho:=[\mathcal{LR}-\mathcal{RL}]\rho=r{\rm Tr}(\rho)\mathcal{L}(\rho_{\delta}),$ (28)
where ${\rm Tr}(\mathcal{L}(\rho))=0$ is used. Hence, a large $n$ is not required when $t_{s}\ll 1$, $rt_{s}\ll 1$ or $\rho_{\delta}$ is close to $\rho_{\rm ss}$. A small $t_{s}$ with moderate $r$ offers a practical regime that may yield substantial acceleration and simultaneously reduce the required number of Trotter steps.
@@
which could be interpreted as a partial and local reset channel (only the $z$-axis is affected, and only local one-body jump operators are involved). As shown in Fig. 3, the relaxation is accelerated significantly.
Figure 3: Dephasing noise induces acceleration of relaxation in a 5-site TFIM. Parameters: (a) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.2\tau_{2}$ (b) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.8\tau_{2}$.